OK I have another one, 3.2+x over 2.5=4.6 subtract 3.2 from each side so I get, x over 2.5=1.4 so then I multiply 2.5 on each side to isolate the x so the answer should be x=3.5 Does this look correct ti you? If so thanks again, if not please explain. Thank you for all your help, Scott Answer by stanbon(75887) 
 3.2+x over 2.5=4.6 subtract 3.2 from each side so I get, x over 2.5=1.4 so then I multiply 2.5 on each side to isolate the x so the answer should be x=3.5
---------------------
3.2 + (x/2.5) = 4.6
---
x/2.5 = 1.4
---
x = 2.5*1.4
---
x = 3.5
============
Cheers,
Stan H.